Explain how you would prepare 10 mL of the Laemli Sample Loading Buffer as given below, using the following stock solutions: 1 M DTT, 10% SDS, 1M Tris HCl pH 6.9, 100% Glycerol, 0.2% stock of Bromophenol Blue.
Answer #1

Answer

Preparation of  10 mL of the 1 X Laemli Sample Loading Buffer

(a) SDS

Stock Concentration (M1) : 10 %

Working Concentration (M2) : 4%

Final volume (V2) : 10ml

Initial Volume: V1

According to molarity formulae

M1V1 = M2V2

10 X V1 = 4 X 10

V1 = 40 /10 = 4 ml

(b) Tris HCl pH 6.9

Stock Concentration (M1) : 1 M

Working Concentration (M2) : 0.125M

Final volume (V2) : 10ml

Initial Volume: V1

According to molarity formulae

M1V1 = M2V2

1 X V1 = 0.125 X 10

V1 = 1.25 ml

(c) Glycerol

Stock Concentration (M1) : 100 %

Working Concentration (M2) : 20 %

Final volume (V2) : 10ml

Initial Volume: V1

According to molarity formulae

M1V1 = M2V2

100 X V1 =20 X 10

V1 = 2ml

d)  Bromophenol Blue

Stock Concentration (M1) : 0.2 %

Working Concentration (M2) : 0.004%

Final volume (V2) : 10ml

Initial Volume: V1

According to molarity formulae

M1V1 = M2V2

0.2 X V1 = 0.004 X 10

V1 = 0.2ml

e) DTT

Stock Concentration (M1) : 1 M

Working Concentration (M2) : 50mM

Final volume (V2) : 10ml

Initial Volume: V1

According to molarity formulae

M1V1 = M2V2

1 X V1 = 50 X 10

V1 = 500 / 1000 ( 1M = 1000mM)

V1 = 0.5ml

So,  

  • SDS : 4ml
  • DTT : 0.5ml
  • Bromophenol Blue: 0.2ml
  • Glycerol : 2ml
  • Tris HCl pH 6.9 : 1.25ml

Total volume : 10 - 7.95= 2.05 ml (Water)